a baloon of mass 100kg is at 100mts from the ground . a man of 50kg mass climbs the rope of the baloon with velocity 0.3m/s relative to the


baloon. in which direction will the baloon move and with what velocity?

Grade:10

1 Answers

vikas askiitian expert
13 years ago

take downward as negative & upward as positive ....here weight of man is balanced by baloon so no net force acts in the system of baloon + man...

due to motion of man in upward direction the baloon will move in downward direction according to momentam conservation...

velocity of man = Vm wrt to baloon

let velocity of baloon is Vb then

from conservation of momentam in Y direction

m(Vm - Vb) - MVb = 0                     ( here momentam is wrt to earth)

mVm = (M+m)Vb                                    ( mass of man & baloon are m & M respectively)

Vb = mVm/(M+m) .......

Vb = 50*0.3/150 = 0.1m/s